# RDA-INR: Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations

### v2

@@ Appendix D Implementation details
D.1 Neural network architectures
Figure D.1: The architectures for the template neural network $f_{\theta}$ and the stationary velocity vector fields $v_{\varphi_{k}}$ in Equation (17). The small red boxes correspond to the latent code input $z$ and the spatial input $x$. Furthermore, the small blue box corresponds to an output, while the rectangles are linear layers with $d_{vel}$ or $d_{\mu}$ output dimensions. Finally, the $\oplus$ and $\otimes$ stand for elementwise addition and scalar/elementwise multiplication, respectively.
Figure D.1 shows the used neural network architectures. For the template neural network, we use nearly the same architecture as the template neural network in [38]. The only difference is that we use ReLU activation functions instead of sine activation functions. In our experiments, we clamp the output of the template neural network to the interval $[-0.5,0.5]$. The stationary velocity vector field neural networks $v_{\varphi_{k}}$ in Equation (17) use the architecture in [21]. However, there are some differences. First, we add an extra linear layer at the output. We also add a scalar multiplication with the function $h_{\epsilon}$. This component ensures the velocity vector field is zero outside $\Omega$, which is needed to let the ODE be a diffeomorphism on $\Omega$.
@@ D.2 Pointwise loss
$L_{pw}=\sum_{t\in T}\sum_{i=1}^{N}\sum_{j=1}^{M}\mathcal{L}_{0.25}\left(\lVert\phi_{i}(p_{ij},t)-p_{j}\rVert_{2}\right),$
where $\mathcal{L}_{0.25}$ is the Huber loss with parameter equal to $0.25$, $T$ is a set of predefined time instances at which to evaluate the pointwise loss, $\{p_{ij}\}_{j=1}^{M}$ are $M$ points from the domain, and $\phi_{i}(\cdot,t):=\phi_{t}^{z_{i}}(\cdot)$ with $\frac{\partial}{\partial t}\phi_{t}^{z}(x)=v_{\varphi}(\phi_{t}^{z}(x),t,z)$ and $\phi_{0}^{z}(x)=x$. For our purposes, we follow [39] and [21] and choose $T=\{\lfloor K/4\rfloor\cdot i\mid 1\leq i\leq K/\lfloor 0.25K\rfloor,i\in\mathbb{N}\}$ with $K$ the number of stationary velocity vector fields in $v_{\varphi}(x,t,z)=\sum_{k=1}^{K}\chi_{[(k-1)/K,k/K)}(t)\cdot v_{\varphi_{k}}(x,z)$.
In [21], the model uses the above loss to learn a template shape with similar features to the training shapes. Hence, similar to our Riemannian regularization $\int_{0}^{1}\lVert v(\cdot,t)\rVert_{V}^{2}\mathrm{d}t$, the pointwise loss is aimed at learning a proper template. Furthermore, note that when choosing $\lVert\cdot\rVert_{V}=\lVert\cdot\rVert_{L_{2}(\Omega)}$ and defining $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=v_{\varphi}(x(t),t,z)$, our Riemannian regularization penalizes large differences between $x(1)$ and $x(0)$. As the pointwise loss penalizes a large $\lVert x(1)-x(0)\rVert_{2}$ as well, the pointwise loss can be seen as a non-Riemannian version of our Riemannian LDDMM regularization.
D.3 Datasets
@@
The synthetic rectangles dataset is created by generating random boxes with edge lengths uniformly distributed in $[0.15,0.85]$. Subsequently, these boxes are rotated using a random rotation matrix. For training the point cloud-based model, we sample 100 000 uniform points and corresponding normals from the meshes. For training using the occupancy data (see Appendix B), we uniformly sample 100 000 points from $\Omega=[-1,1]^{3}$ and calculate the signed distance to the boxes. Subsequently, we calculate the occupancy values from these signed distance values. The training dataset consists of 100 randomly generated parallelograms, while the test dataset consists of 20 parallelograms.
@@ D.4 Training details
For the experiment done in Appendix B with the occupancy data, we first sample 5000 points in $\Omega$ with their accompanying occupancy value. Half of the points lie inside, while the other half lie outside the shape. We calculate the binary cross entropy loss by averaging the binary cross entropies between the ground truth occupancy value and the estimated occupancy probability.
We update the neural network parameters and the latent codes $z_{i}$ via separate Adam optimizers for the latent codes, the template NN $f_{\theta}$, and velocity field NN $v_{\varphi}$. The initial learning rate for the latent codes is $10^{-3}$, for the $f_{\theta}$ parameters $5\cdot 10^{-4}$, and for the $v_{\varphi}$ parameters $5\cdot 10^{-4}$. Moreover, for each Adam optimizer, we use a learning rate scheduler that multiplies the learning rate with $0.7$ every 250 epochs. Finally, we bound the latent codes $z_{i}$ to be within the unit sphere. For the remainder of the hyperparameters, see Table D.1.
@@ D.5 Inference details
$\min_{z}\mathcal{D}_{rec}((\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta},\mathcal{S})+\gamma\lVert z\rVert_{2}^{2},$
@@
Table D.1: The hyperparameter values every trained model uses on a specific dataset (rectangles or liver). Here $N_{h}$ is the number of linear layers between the first and last linear layer in the template neural network (as presented in Figure D.1). Furthermore, $c_{\text{PW}}$ denotes the regularization constant for the pointwise loss when training the model that uses this loss as the deformation regularizer.
@@
D.6 DAE training and inference details
@@
First, we used Deformetrica’s atlas estimation [87] to compute an initial template. We set the noise standard deviation of the current similarity metric to 0.01 and the kernel width to $0.1$. For the deformation kernel width, we used a value of $0.15$. Furthermore, we used ScipyLBFGS as the optimizer with an initial step size of $3\cdot 10^{-5}$ and a very low convergence tolerance. Once we obtained the template, we employed Deformetrica’s PGA function to obtain a $32$-dimensional latent space. We initialized the initial template with the estimated template and ran PGA with the same parameters as for the atlas estimation.
We utilized the calculated PGA to initialize the DAE model. We use $15000$ epochs for this initialization. After that, we trained the DAE model for $15000$ epochs with a learning rate of $5\cdot 10^{-4}$ and a batch size of $32$. We kept the template fixed as learning it resulted in a lot of instability. We chose a splatting grid size of $16$, a deformation grid size of $32$, and a kernel width of $0.15$ for the splatting layer and the current metric. We multiplied the data-fitting term by 1e4 and the velocity field regularization term by $0.1$.
To evaluate the model, we used the encoder to estimate a latent vector. For the test dataset, we also performed an additional optimization step for $1500$ iterations on the latent code using an Adam optimizer with a learning rate of $5\cdot 10^{-3}$. This additional optimization step corresponds to the DAE+ version of the model, which was used to improve reconstruction performance on the test set.

### v3

@@ Appendix D Implementation details
D.1 Parameterizing and solving the ordinary differential equation
(20) $v_{\varphi}(x,t,z)=\sum_{k=1}^{K}\chi_{[\frac{k-1}{K},\frac{k}{K})}(t)\cdot v_{\varphi_{k}}(x,z).$
@@ D.2 Pointwise loss
$L_{pw}=\sum_{t\in T}\sum_{i=1}^{N}\sum_{j=1}^{M}\mathcal{L}_{0.25}\left(\left\lVert\phi_{i}(p_{ij},t)-p_{ij}\right\rVert_{2}\right),$
where $\mathcal{L}_{0.25}$ is the Huber loss with parameter equal to $0.25$, $T$ is a set of predefined time instances at which to evaluate the pointwise loss, $\{p_{ij}\}_{j=1}^{M}$ are $M$ points from the domain, and $\phi_{i}(\cdot,t):=\phi_{t}^{z_{i}}(\cdot)$ with $\frac{\partial}{\partial t}\phi_{t}^{z}(x)=v_{\varphi}(\phi_{t}^{z}(x),t,z)$ and $\phi_{0}^{z}(x)=x$. The Huber loss is given by:
$L_{\delta}(a):=\begin{cases}\frac{1}{2}a^{2}&\mathrm{if}|a|\leq\delta\\ \delta\cdot\left(|a|-\frac{1}{2}\delta\right)&\mathrm{otherwise}\end{cases}$
We follow [85] and [63] and choose $T=\{\lfloor K/4\rfloor\cdot i\mid 1\leq i\leq K/\lfloor 0.25K\rfloor,i\in\mathbb{N}\}$ with $K$ the number of stationary velocity vector fields in $v_{\varphi}(x,t,z)=\sum_{k=1}^{K}\chi_{[(k-1)/K,k/K)}(t)\cdot v_{\varphi_{k}}(x,z)$.
In [63], the pointwise loss is used to learn a template shape with similar features as the training shapes, which is one goal of our Riemannian regularization $\int_{0}^{1}\left\lVert v(\cdot,t)\right\rVert_{V}^{2}\mathrm{d}t$ as well. Furthermore, note that when choosing $\left\lVert\cdot\right\rVert_{V}=\left\lVert\cdot\right\rVert_{L_{2}(\Omega)}$ and defining $\frac{\mathrm{d}}{\mathrm{d}t}x(t)=v_{\varphi}(x(t),t,z)$, our Riemannian regularization penalizes large differences between $x(1)$ and $x(0)$. As the pointwise loss also penalizes a large $\left\lVert x(1)-x(0)\right\rVert_{2}$, it can be seen as a non-Riemannian version of our Riemannian LDDMM regularization.
D.3 Neural network architectures
Figure 15 shows the used neural network architectures. For the template neural network, we use nearly the same architecture as the template neural network in [25]. The only difference is that we use ReLU activation functions instead of sine activation functions. In our experiments, we clamp the output of the template neural network to the interval $[-0.5,0.5]$. The stationary velocity vector field neural networks $v_{\varphi_{k}}$ in Equation (20) use the architecture in [63]. However, there are some differences. First, we add an extra linear layer at the output. We also add a scalar multiplication with the function $h_{\epsilon}$. This component ensures the velocity vector field is zero outside $\Omega$, which is needed to let the ODE be a diffeomorphism on $\Omega$.
Figure 15: Neural network architectures. Architectures for the template neural network $f_{\theta}$ and the stationary velocity vector fields $v_{\varphi_{k}}$ in Equation (20). The small red boxes indicate the latent code input $z$ and the spatial input $x$. Furthermore, the small blue box is an output, while the rectangles are linear layers with $d_{vel}$ or $d_{\mu}$ output dimensions. Finally, the $\oplus$ and $\otimes$ stand for elementwise addition and scalar/elementwise multiplication, respectively.
D.4 Datasets
@@
The synthetic rectangles dataset is created by generating random boxes with edge lengths uniformly distributed in $[0.15,0.85]$. Subsequently, these boxes are rotated using a random rotation matrix. For training the point cloud-based model, we sample 100 000 uniform points and corresponding normals from the meshes. For training using the occupancy data (see Appendix B.1), we uniformly sample 100 000 points from $\Omega=[-1,1]^{3}$ and calculate the signed distance to the boxes. Subsequently, we calculate the occupancy values from these signed distance values. The training dataset consists of 100 randomly generated parallelograms, while the test dataset consists of 20 parallelograms.
@@ D.5 Training details
For the experiment done in Appendix B.1 with the occupancy data, we first sample 5000 points in $\Omega$ with their accompanying occupancy value. Half of the points lie inside, while the other half lie outside the shape. We calculate the binary cross entropy loss by averaging the binary cross entropies between the ground truth occupancy value and the estimated occupancy probability.
We update the neural network parameters and the latent codes $z_{i}$ via separate Adam optimizers for the latent codes, the template NN $f_{\theta}$, and velocity field NN $v_{\varphi}$. The initial learning rate for the latent codes is $10^{-3}$, for the $f_{\theta}$ parameters $5\cdot 10^{-4}$, and for the $v_{\varphi}$ parameters $5\cdot 10^{-4}$. Moreover, for each Adam optimizer, we use a learning rate scheduler that multiplies the learning rate with $0.7$ every 250 epochs. Finally, we bound the latent codes $z_{i}$ to be within the unit sphere. For the remainder of the hyperparameters, see Table 7.
@@ D.6 Inference details
$\min_{z}\mathcal{D}_{rec}((\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta},\mathcal{S})+\gamma\left\lVert z\right\rVert_{2}^{2},$
@@
Table 7: Hyperparameters for model training. The hyperparameter values every trained model uses on a specific dataset (rectangles or liver). Here $N_{h}$ is the number of linear layers between the first and last linear layer in the template neural network (as presented in Figure 15). Furthermore, $c_{\text{PW}}$ denotes the regularization constant for the pointwise loss when training the model that uses this loss as the deformation regularizer.
@@
D.7 DAE training and inference details
@@
First, we compute an initial template via Deformetrica’s atlas estimation [15]. We set the noise standard deviation of the current similarity metric to 0.01 and the kernel width to $0.1$. For the deformation kernel width, we used a value of $0.15$. Furthermore, we used ScipyLBFGS as the optimizer with an initial step size of $3\cdot 10^{-5}$ and a very low convergence tolerance. Once we obtained the template, we employed Deformetrica’s PGA function to obtain a $32$-dimensional latent space. We initialized the initial template with the estimated template and ran PGA with the same parameters as for the atlas estimation.
We utilized the calculated PGA to initialize the DAE model. We use $15000$ epochs for this initialization. After that, we trained the DAE model for $15000$ epochs with a learning rate of $5\cdot 10^{-4}$ and a batch size of $32$. We kept the template fixed as learning it resulted in a lot of instability. We chose a splatting grid size of $16$, a deformation grid size of $32$, and a kernel width of $0.15$ for the splatting layer and the current metric. We multiplied the data-fitting term by $10^{4}$ and the velocity field regularization term by $0.1$.
To evaluate the model, we used the encoder to estimate a latent vector. For the test data, we performed an additional optimization step for $1500$ iterations on the latent code using an Adam optimizer with a learning rate of $5\cdot 10^{-3}$. This optimization step is the DAE+ version of the model, which was used to improve reconstruction performance on the test set.